# CER: Confidence Enhanced Reasoning in LLMs

### v1

@@ Abstract
Ensuring the reliability of Large Language Models (LLMs) in complex reasoning tasks remains a formidable challenge, particularly in scenarios that demand precise mathematical calculations and knowledge-intensive open-domain generation. In this work, we introduce an uncertainty-aware framework designed to enhance the accuracy of LLM responses by systematically incorporating model confidence at critical decision points. We propose an approach that encourages multi-step reasoning in LLMs and quantify the confidence of intermediate answers such as numerical results in mathematical reasoning and proper nouns in open-domain generation. Then, the overall confidence of each reasoning chain is evaluated based on confidence of these critical intermediate steps. Finally, we aggregate the answer of generated response paths in a way that reflects the reliability of each generated content (as opposed to self-consistency in which each generated chain contributes equally to majority voting). We conducted extensive experiments in five datasets, three mathematical datasets and two open-domain datasets, using four LLMs. The results consistently validate the effectiveness of our novel confidence-aggregation method, leading to an accuracy improvement of up to 7.4% and 5.8% over baseline approaches in math and open-domain generation tasks, respectively. [^1] Code is publicly available at https://github.com/Aquasar11/CER.
@@ 1 Introduction
Figure 1: Illustration of Confidence-Enhanced Reasoning (CER) in LLMs. On the left, we demonstrate the CER framework. Given an input query, the LLM generates three independent outputs using temperature sampling ($T=1$). Intermediate answers are bolded, and final answers are highlighted. The confidence of each output is computed, and the most weighted-confident answer—125—is selected. On the right, we illustrate the confidence calculation for the first output. We use multiplication as the step-wise aggregator function ($f$) and weighted averaging ($wa$) as the path-wise aggregator function ($g$). Since the answer 125 appears in both step 4 and the final answer, we mark its first occurrence with * for clarity. The full question and responses from the LLM are provided in Appendix F.
@@ 3.2 Method
MATH 20.4 13.6 19.0 20.2 20.0 19.6 17.0 18.0 (-2.4%)
MathQA 20.8 12.4 22.6 20.0 19.4 22.6 20.2 22.6 (+0%)
Average 34.4 24.2 32.4 33.0 33.1 32.6 27.3 35.2 (+0.8%)
@@ 4.1 Experimental Setup
More details on the formulation and the aggregation approach of confidence-based methods are provided in the Appendix G.
@@ 4.3 Ablation Studies
5 Conclusion
In this paper, we introduced a lightweight framework that enhances performance on various reasoning tasks by relying solely on the model’s output logits without the need for fine-tuning or task-specific prompts. Our approach bridges the gap between reasoning and uncertainty estimation in LLMs. Through extensive experiments, we validated our proposed functions and demonstrated the CER algorithm’s effectiveness as a general performance enhancement framework. Our findings show that the framework is robust across different model sizes and architectures. In this study, we focus only on numerical outputs in mathematical reasoning. However, with minor modifications, our approach can also handle non-numerical outputs, such as mathematical proofs. Future research could further extend our framework to mathematical reasoning tasks with non-numerical final answers.
Limitations
Our work has several notable limitations. First, the framework has been applied only to a narrow range of tasks, specifically those involving mathematical reasoning and knowledge-intensive questions. Furthermore, our approach relies on access to the model output logits; therefore, our method is not applicable in scenarios where these logits are unavailable.
@@ Appendix D More Results
Appendix E Further Exploration of the Dataset
@@
Appendix F Full question and responses related to the main figure
@@
Appendix G Examination of Confidence Baselines

### v2

@@ Abstract
Ensuring the reliability of Large Language Models (LLMs) in complex reasoning tasks remains a formidable challenge, particularly in scenarios that demand precise mathematical calculations and knowledge-intensive open-domain generation. In this work, we introduce an uncertainty-aware framework designed to enhance the accuracy of LLM responses by systematically incorporating model confidence at critical decision points. We propose an approach that encourages multi-step reasoning in LLMs and quantify the confidence of intermediate answers such as numerical results in mathematical reasoning and proper nouns in open-domain generation. Then, the overall confidence of each reasoning chain is evaluated based on confidence of these critical intermediate steps. Finally, we aggregate the answer of generated response paths in a way that reflects the reliability of each generated content (as opposed to self-consistency in which each generated chain contributes equally to majority voting). We conducted extensive experiments in five datasets, three mathematical datasets and two open-domain datasets, using four LLMs. The results consistently validate the effectiveness of our novel confidence-aggregation method, leading to an accuracy improvement of up to 7.4% and 5.8% over baseline approaches in math and open-domain generation tasks, respectively. [^1] Code is publicly available at https://github.com/sharif-ml-lab/CER.
@@ 1 Introduction
Figure 1: Illustration of Confidence-Enhanced Reasoning (CER) in LLMs. On the left, we demonstrate the CER framework. Given an input query, the LLM generates three independent outputs using temperature sampling ($T=1$). Intermediate answers are bolded, and final answers are highlighted. The confidence of each output is computed, and the most weighted-confident answer—125—is selected. On the right, we illustrate the confidence calculation for the first output. We use multiplication as the step-wise aggregator function ($f$) and weighted averaging ($wa$) as the path-wise aggregator function ($g$). Since the answer 125 appears in both step 4 and the final answer, we mark its first occurrence with * for clarity. The full question and responses from the LLM are provided in Appendix G.
@@ 3.2 Method
MATH 20.4 13.6 19.0 20.2 20.0 19.6 17.0 24.0 (+3.6%)
MathQA 20.8 12.4 22.6 20.0 19.4 22.6 20.2 22.6 (+0.0%)
Average 34.4 24.2 32.4 33.0 33.1 32.6 27.3 37.2 (+2.8%)
@@ 4.1 Experimental Setup
More details on the formulation and the aggregation approach of confidence-based methods are provided in the Appendix H.
@@ 4.3 Ablation Studies
Recent Reasoning-based Models
To evaluate appling our method on recent reasoning models, DeepSeek-R1-Distill-Qwen-7B model Guo et al. (2025), a distilled variant of the state-of-the-art DeepSeek-R1 architecture, is utilized. This model is specifically optimized for mathematical reasoning tasks using the GRPO reinforcement learning method. We compare employing our method against several baseline methods on two benchmarks: GSM8K, which focuses on mathematical reasoning, and TriviaQA, which targets open-domain question answering. The obtained results are shown in Table 3.
Models & Datasets Self-Consistency P(True) PE NL NE LL Greedy CER
GSM8K 89.8 82.6 83.8 85.2 83.8 84.0 87.2 90.2
Trivia QA 22.4 15.2 16.0 15.4 14.8 15.8 19.4 24.4
Table 3: Comparison of CER with baselines on DeepSeek-R1-Distill-Qwen-7B across the GSM8K and TriviaQA datasets.
5 Discussion and Conclusion
In this paper, we introduced a lightweight framework that enhances performance on various reasoning tasks by relying solely on the model’s output logits without the need for fine-tuning or task-specific prompts. Our approach bridges the gap between reasoning and uncertainty estimation in LLMs. Through extensive empirical analysis, we observe some key findings: (1) Aggregating multiple responses based on the model’s confidence leads to more reliable outcomes. Unlike self-consistency methods, which generate multiple reasoning paths but assign them equal weight, CER employs a confidence-based approach that assigns weights to each path and aggregates them accordingly. (2) Considering intermediate answer tokens to assess confidence is more effective than using all tokens in the chain. We observe that at the end of each step, the model is expected to arrive at a certain level of confidence in its output, while some degree of uncertainty is natural throughout a thought due to an incomplete or evolving reasoning step. Therefore, a CoT should not be penalized for these natural uncertainties throughout its thoughts. As a result, the overall undesired uncertainty in the reasoning chain is inferred by analyzing the confidence levels of only the critical tokens that constitute the intermediate and final answers, while discarding the confidence of other tokens that may lead to misleading conclusions. To achieve this, CER encourages the LLM to express intermediate answers through reasoning and leverages these tokens—like numerical values in mathematical reasoning and proper nouns in open-domain generation. (3) Assigning higher weights to reasoning steps closer to the final answer improves performance. By combining these findings, CER thus enhances the reasoning capabilities of large language models.
6 Limitations
Our work has several notable limitations. First, the framework has currently been applied only to a limited range of tasks, specifically those that involve mathematical reasoning and knowledge-intensive questions. This restriction arises from the nature of our method, which requires intermediate answers or clear indicators of intermediate results for operation. Additionally, our approach depends on the access to the model output logits. Consequently, our method cannot be employed in scenarios where logits are inaccessible, such as with black-box settings. However, it can be adapted to these settings by modifying the step-wise aggregation, prompting the LLM for intermediate answers, estimating confidence from verbalized confidence at each step, and applying path-wise aggregation to compute overall confidence. Also, in this study, we focus only on numerical outputs in mathematical reasoning. However, with minor modifications, our approach can also handle non-numerical outputs, such as mathematical proofs. Future research could further extend our framework to mathematical reasoning tasks with non-numerical final answers.
Acknowledgments
We would like to thank everyone who supported us throughout the development of this paper. Special thanks to Dr. Kazemi and Dr. Hassanifard at MCINext for their generous support and for providing the GPU and server resources essential for our research.
@@ Appendix D More Results
Appendix E Additional Prompt Formats and Evaluation
We evaluated our method on all models using two additional prompt formats on the GSM8K data set. The second format closely resembles our original prompt but replaces Answer:[answer] with Response:[response] for intermediate steps. The third format offers the most flexibility, allowing the LLM to generate intermediate answers freely at the end of each step without enforcing a specific structure. However, to ensure that the final answer remains easily identifiable and error-free, we still require the model to provide it in a predefined format.
E.1 Prompt Formats
The second prompt, which offers greater flexibility than the original, is presented in Figure 6, while the third prompt, the most flexible of the three, is shown in Figure 7.
**Objective** Carefully work through the problem step by step. For each step, perform any required reasoning, and express the response at the end of the step. Your response for each intermediate step should be in the format Response: [response]. After completing the steps, provide the final response based on the reasoning developed throughout the process. **Important Rules** Your response should end with: The final response is [response], where [response] is the final response to the problem. Q: <question>
Figure 6: Second Prompt (more flexible)
**Objective** Carefully work through the problem step by step. For each step, perform any required reasoning and express the answer at the end of the step. After completing the steps, provide the final answer based on the reasoning developed throughout the process. **Important Rules** Your response should end with The final answer is [answer], where [answer] is the final response to the problem. Q: <question>
Figure 7: Third Prompt (most flexible)
E.2 Prompt Results
Table 6 presents the results of different models evaluated on the GSM8K dataset using these prompts, while Table 7 displays the comparison results of our primary models on the MATH and Allen AI Math datasets on the primary prompt and the most flexible one.
Model Original Prompt Second Prompt Third Prompt
Llama 3.1 8B 90.0 88.0 90.6
Mistral 7B 65.2 62.6 67.8
Llama 3.2 3B 82.6 81.2 80.4
Olmo 2 7B 88.8 88.4 90.2
Table 6: Accuracy on GSM8K with different prompt formats
MATH Dataset Allen AI Math Dataset
Model Original Prompt Third Prompt Original Prompt Third Prompt
Llama 3.1 8B 58.2 57.4 68.2 66.0
Mistral 7B 18.0 18.6 22.6 21.8
Table 7: A comparison of accuracy on the MATH and Allen AI Math datasets using the main prompt (Figure 4) and the most flexible prompt (Figure 7).
Appendix F Further Exploration of the Dataset
@@
Appendix G Full question and responses related to the main figure
@@
Appendix H Examination of Confidence Baselines